OpenStudy (anonymous):

Solve the equation by taking square roots: 5x^2-3x+2=4x^2-3x+2

OpenStudy (anonymous):

The answer on the answer sheet is 0

OpenStudy (anonymous):

x = 0

OpenStudy (anonymous):

Show work

OpenStudy (anonymous):

Simplify x (5 x-3)+2 = x (4 x-3)+2 x^2 = 0 x = 0

OpenStudy (anonymous):

x^2 -3x + 3x + 2 - 2 = 0 x^2 = 0 x = 0

OpenStudy (anonymous):

How do I go from: x(5x-3)+2=x(4x-3)+2

OpenStudy (anonymous):

expand the the equation on both sides then cancel out so\[5x^2-3x+3=4x^2-3x+2\] then take away all like terms \[5x^2-3x=4x^2-3x\] \[x^2=0\]

OpenStudy (anonymous):

should be +2 sorry on the LHS